OpenStudy (anonymous):

An open box is made from a square piece of sheet metal 19 inches on a side by cutting identical squares from the corners and turning up the sides. Express the volume of the box, V, as a function of the length of the side of the square cut from each corner, x. A. V(x) = 361x B. V(x) = (19 - 2x)2 C. V(x) = x(19 - 2x) D. V(x) = x(19 - 2x)2

OpenStudy (campbell_st):

draw diagram  the dimensions of the area that will be used as the base will be length = (19 -2x) so the area of the base of the box will be \[A = (19 -2x)^2\] the height of the sides of the box will be x ( the dimension of the cut out ) Volume = Base area * height so \[ V = (19 -x)^2 \times x\] or \[V = x(19 -2x)^2\]